OpenStudy (anonymous):

solve –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (zepp):

First of all, what can you do with the left side?

OpenStudy (anonymous):

well I need to distribute the -2 so i have -2x-3 is greater than or equal to 5-(x+3)?

OpenStudy (zepp):

Nope, look: –2(x – 3) -2*(x)-(-2)*3 -2x-(-6) -2x+6 It should be like this, then you have (-2x+6) ≥ 5 – (x + 3)

OpenStudy (zepp):

You could pass the (x+3) to the left side by adding (x+3) to both side (-2x+6) ≥ 5 – (x + 3) (-2x+6) + (x+3) ≥ 5

OpenStudy (anonymous):

I am so confused on how to do this so you multiplied the -2 then add the (x+3)? why? how would I get to the conclusion of this?

OpenStudy (anonymous):

-2(x – 3) ≥ 5 – (x + 3) -2*(x)-(-2)*3 -2x-(-6) -2x+6 (-2x+6) ≥ 5 – (x + 3) (-2x+6) + (x+3) ≥ 5

OpenStudy (anonymous):

-2(x – 3) ≥ 5 – (x + 3) -2x+6>= 5-x-3 -X>=-4 x<=4

OpenStudy (anonymous):

were u solving for x lol

OpenStudy (ash2326):

@Supernalstudier did you understand?

OpenStudy (anonymous):

yes I was solving for x did you multiply the -2 to the -3? and distribute the 5?

OpenStudy (anonymous):

yes multiply the -2 to -3 to expand the bracket

OpenStudy (anonymous):

jus solve as though the >= sign was = you only change its direction when u multiply both sides with a minus sign

OpenStudy (ash2326):

–2(x – 3) ≥ 5 – (x + 3) Let's begin step by step First Let's evaluate the left hand side \[-2(x-3)\] Multiply 2 to both the terms inside the bracket \[-(2x-6)\] Now as we have "-" outside, change the sign of the terms inside the bracket, we get \[-2x+6\] Did you understand till here?

OpenStudy (anonymous):

Thank you all so much I really need help with inequalities I defiantly understand a little bit better and am going to try the next problem on my own

OpenStudy (ash2326):

Great @Supernalstudier

OpenStudy (anonymous):

one question though why did the -2 not cancel out?

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) -2x+6>= 5-x-3 -X>=-4 x<=4 if you multiplied it to both sides wouldn't it cancel out?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

why would u want to multiply -2 to both sides

OpenStudy (anonymous):

you only multiply it to the -3? does that mean I did this one wrong? 1. x – 3 > –9 +3 +3 x>-6

OpenStudy (anonymous):

when u solve problems like that do not add to both sides x-3>-9 x>-9 +3 x>-6

OpenStudy (anonymous):

instead just bring them over

OpenStudy (anonymous):

and when u bring them over to the other side the sign changes

OpenStudy (ash2326):

@Supernalstudier you did right, We will add 3 to both sides

OpenStudy (ash2326):

But here \[-2(x-3)\ge 5-(x+3)\] As we already have -2 in the inequality, we will multiply it to only the left hand side

OpenStudy (ash2326):

This is like \[2\times(x)=4\] here we already have 2 on the left hand side But I can still add, sub, multiply or divide any no. to both sides. We use these both sides operations to get just the variable no number along with it example here I Have \[2\times (x)=4\] I have a 2 with x. I don't like it. I need to remove it I'll divide both sides by 2 and then I'll get \[x=2\] So it's just x. So we always try this in any equation or inequality.

OpenStudy (ash2326):

Did you get this?

OpenStudy (anonymous):

I understand if you have 2x (x) =4 x=4 that one is easier but I would use the same concept for the more complex equations?

OpenStudy (ash2326):

Yeah it's applicable everwhere

OpenStudy (anonymous):

thank you